OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 1) and (3, 4)

OpenStudy (anonymous):

and Input in standard form the equation of the given line. The line with m = -1 /2 and b = 1

OpenStudy (anonymous):

for the equation of line passing through points (x1,y1) and (x2,y2) is y-y1 = m*(x-x1) where m = slope \[m = \frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

I have that! I just need the answer because I don't have it right

OpenStudy (anonymous):

show us your steps it will help us to find the error

OpenStudy (anonymous):

4-1=3 (which is y2-y1) 3-1=2 (which is x2-x1) I get 3/2 But to put it in ax+by=c isnt working

OpenStudy (anonymous):

put it in y-y1 = m*(x-x1) then rearrange to this general form Ax+By =C